03*. A truck travelling at 22.5 m/s decelerates at 2.27 m/s2.


(a) How much time does it take for the truck to stop? [9.91s]


(b) How far does it travel while stopping? [112 m]


(c) How far does it travel during the third second after the brakes are applied? [16.8 m]

1
Expert's answer
2022-06-09T09:59:22-0400

Given:

"v_0=22.5\:\rm m/s"

"a=-2.27\:\rm m/s^2"


(a)

"v_f=v_0+at=0"

"t=-\frac{v_0}{a}=-\frac{22.5}{-2.27}=9.91\:\rm s"

(b)

"d=v_0t+at^2/2\\
=22.5*9.91+(-2.27)*9.91^2/2=112\:\rm m"

(c)

"\Delta d=d(3)-d(2)\\
=(22.5*3+(-2.27)*3^2/2)\\
-(22.5*2+(-2.27)*2^2/2)=16.8\:\rm m"